OpenStudy (anonymous):

Coughing forces the trachea to contract, which affects the velocity v of the air passing through the trachea. Suppose the velocity of the air during coughing is v = k(R-r)r2 where k and R are constants, R is the normal radius of the trachea, and r is the radius during coughing. What radius will produce the maximum air velocity? @Calculus1

OpenStudy (anonymous):

v = k(R - r)r^2 - is this the formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[v=k(R-r)r^2\]\[v=kRr^2-kr^3\]\[v'=2kRr-2kr^2=r(2kR-2kr)=0\]since r can't be zero,\[2kR-2kr=0\]\[r=\frac{2kR}{2k}=R\]thus a radius of R will produce the max air velocity

OpenStudy (anonymous):

wait, miscalc

OpenStudy (anonymous):

1/2R, 2/3R, 3/2R, or 4/9 R are choices

OpenStudy (anonymous):

\[r=\frac{2}{3}R\]***

OpenStudy (anonymous):

howd you get that mahone?

OpenStudy (anonymous):

dv/dr = 2kRr - 3kr^2 = 0 r (2kR - 3kr) = 0 r = 2R / 3

OpenStudy (anonymous):

\[v=k(R-r)r^2\]\[v=kRr^2-kr^3\]\[v'=2kRr-3kr^2=r(2kR-3kr)=0\]\[r=\frac{2}{3}R\]

OpenStudy (anonymous):

just tell your teacher that \[\lim r\to0=\infty\]infinite air velocity ftw

OpenStudy (anonymous):

mahone i dont care what anyone says... u are the man. You too jimmy

OpenStudy (anonymous):

lol 0- ty treyhud